OpenStudy (anonymous):

Let \(f (x) = \sin (11 x)\). What is the fourth derivative of \(f (x)\)? \(\frac{d^4}{d x^4} \sin (11 x) =?

OpenStudy (anonymous):

11^4 sin(11x)

OpenStudy (anonymous):

kindly explain it too

OpenStudy (anonymous):

\[11^4 \sin(11x)\]

OpenStudy (anonymous):

explain

OpenStudy (anonymous):

what is \[\frac{d}{dx}\sin(ax)\]equal to?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

o yes acos(ax) right?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

so what is\[\frac{d}{dx}a\cos(ax)\]equal to?

OpenStudy (anonymous):

-a^2sin(ax)?

OpenStudy (anonymous):

yes :-D what is \[\frac{d}{dx}-a^2\sin(ax)\]equal to?

OpenStudy (anonymous):

ok ok i get it thanks :)